how do you write a let statement and equation for this question?
the smaller of two numbers is 1/5 of the larger number. if the sum of the two numbers is 90, find both numbers.

Answer by checkley77(12844)

x+x/5=90
(5x+x)/5=90
6x=90*5
6x=450
x=450/6
x=75 ans for the larger number.
75/5=15 for the smaller number.
Proof:
75+15=90
90=90